OpenStudy (starlord1):

9 •72 = 9 •____ + 9 • 2 = 648

OpenStudy (maddy1251):

What are you trying to find, exactly?

OpenStudy (maddy1251):

So it is saying 9 times 72 = 9 times ? + 9 times 2 = 648?

OpenStudy (starlord1):

yes

OpenStudy (maddy1251):

Well we do know 9 x 72 = 648, we just gotta find I am assuming the factors to find what its stretched out version will equal 648.

OpenStudy (maddy1251):

9 x 2 = 18.

OpenStudy (maddy1251):

It will be 70.

OpenStudy (maddy1251):

I took 648 and subtracted 18 to get 630.

OpenStudy (maddy1251):

From there, I divided by 9.

OpenStudy (maddy1251):

9 x 70 + 9 x 2 = 648.

OpenStudy (maddy1251):

It's basically this ;

OpenStudy (maddy1251):

9x+18=648

OpenStudy (maddy1251):

9x+18-18=648-18

OpenStudy (maddy1251):

9x=630

OpenStudy (maddy1251):

9x/9=630/9 x=70